three students are each asked to select four numbers less than ten and multiply them together. they all have the same answer, even though they only share one number the same. what is the answer and the shared number
Answer by Edwin McCravy(20055)

The answer is 0 and the shared number is 0.

That's because if the all pick a 0 the product will be 0.

For instance: 

the 1st person could pick the numbers 0,1,2,3, and 0×1×2×3 = 0
the 2nd person could pick the numbers 0,4,5,6, and 0×4×5×6 = 0
the 3rd person could pick the numbers 0,7,8,9, and 0×7×8×9 = 0

They all share only one number, 0, and their products are all 0.

Edwin